The following observations were randomly sampled from a Normal distribution. Using this data, construct a 99% confidence interval for the mean and interpret the interval. 8.11 12.06 10.01 12.01 8.59 9.35 10.79 11.74 7.42 11.72 8.44 7.67 10.93 8.42 8.50


1
Expert's answer
2021-06-11T02:37:33-0400
"\bar{x}=\dfrac{\displaystyle\sum_{i=1}^nx_i}{n}=\dfrac{\displaystyle\sum_{i=1}^{15}x_i}{15}"

"=\dfrac{1}{15}(8.11+12.06 +10.01 +12.01+ 8.59"

"+9.35+ 10.79 +11.74+ 7.42+ 11.72"

"+ 8.44+ 7.67+ 10.93+ 8.42+ 8.50=\dfrac{145.76}{15}"

"\approx9.717333"

"s^2=\\dfrac{\\displaystyle\\sum_{i=1}^{15}(x_i-\\bar{x})^2}{15-1}\\approx \t2.843450"

"s=\sqrt{s^2}\approx1.686253"

The critical value for "\alpha=0.01" and "df=n-1=15-1=14" degrees of freedom is "t_c=2.976842."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}-t_c\times\dfrac{s}{\sqrt{n}})"

"=(9.717333-2.976842\times\dfrac{1.686253}{\sqrt{15}},"

"9.717333+2.976842\times\dfrac{1.686253}{\sqrt{15}})"

"\approx(8.421, 11.013)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "8.421<\mu <11.013," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(8.421, 11.013)."